OpenStudy (ria23):

State the Domain and Range of the function 2x^3-1/5x^3-16 Please show me how yhu worked it out~

OpenStudy (isaiah.feynman):

Its a rational function, so the denominator cannot equal zero. The domain is the set of real numbers that make the function defined.

OpenStudy (ria23):

So, I would look at the highest exponent right? That's what I was told... But the highest exponent is 3 and it's on top and bottom.

OpenStudy (mathmath333):

is it this \(\dfrac{2x^3-1}{5x^3-16}\)

OpenStudy (ria23):

Yes

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{5x^3-16\neq 0 }\end{align}\)

OpenStudy (ria23):

That's... The domain? Or the first step I would take to find it.

OpenStudy (mathmath333):

that is not the domain ,the solutions which this equation will give will not be part of the domain

OpenStudy (mathmath333):

here u can exclude the imaginary solutions

OpenStudy (ria23):

Will it show me, like... How it did it? The steps so I know how to solve the other problems like this?

OpenStudy (mathmath333):

u can solve it by formula \(\Large a^3-b^3=(a-b)(a^2+ab+b ^2)\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{5x^3-16=0\\ x^3-\dfrac{16}{5}\\ =x^3-2^3(\dfrac{ 2}{5})^{\frac{3}{3}} }\end{align}\)

OpenStudy (mathmath333):

\(\Large a^3-b^3=(a-b)(a^2+ab+b ^2)\) from this formula \(\large\tt \begin{align} \color{black}{(a^2+ab+b ^2) }\end{align}\) part will give imaginary solutions so u have to just ignore it

OpenStudy (mathmath333):

so we are left with \(\Large\tt \begin{align} \color{black}{(a-b) }\end{align}\)

OpenStudy (mathmath333):

r u getting it?

OpenStudy (mathmath333):

\(\Large\tt \begin{align} \color{black}{(a-b) \implies x-(2\sqrt[3]{\dfrac{ 2}{5}})=0\\ x=(2\sqrt[3]{\dfrac{ 2}{5}}) }\end{align}\)

OpenStudy (mathmath333):

domain = \(\Large\tt \begin{align} \color{black}{x\in \mathbb{R}:x\neq (2\sqrt[3]{\dfrac{ 2}{5}}) }\end{align}\) where \(\Large \mathbb{R}\)=real numbers

OpenStudy (ria23):

I... Kind of get it... So then the range is... what the y values can be. which would be... \[r(x)\in \mathbb{R},0>x>\infty \] ?

OpenStudy (mathmath333):

here when \(\large\tt \begin{align} \color{black}{f(x)=\dfrac{2}{5}\\ \dfrac{2}{5}=\dfrac{2x^3-1}{5x^3-16}\\ 10x^3-32=10x^3-5\\}\end{align}\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ -32\neq -5\\}\end{align}\)

OpenStudy (ria23):

So how would I write that in terms of range?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{f(x)\in \mathbb{R}:f(x) \neq \dfrac{2}{5} }\end{align}\)

OpenStudy (ria23):

>.< I was way wrong. If I do another one and try it, do yhu mind if I tag yhu and yhu can check what I did? I have 2 more kind of like this.

OpenStudy (mathmath333):

i will try

OpenStudy (ria23):

Thank yhu~ c: